OpenStudy (anonymous):

factor the algebric expression x2-z4

OpenStudy (anonymous):

Use \[ a^2 - b^2=(a-b)(a+b) \]

OpenStudy (anonymous):

with \[ a=x\\ b=z^2 \]

OpenStudy (anonymous):

\[ x^2-z^4=(x-z^2)(x+z^2) \]